Assume that you have four different processes with the following attributes:

Process   Arrival time.   CPU Burst.   I/O Burst Total CPU time

A. 0 4 4 9

B 3 2 3 7

C 6 5 1 11

D 12 1 1 5

As we did in class, perform a scheduling simulation using these four processes according to the following algorithms:

1) First Come First Serve

2) Round Robin with time slice = 1

3) Round Robin with time slice = 3

4) Shortest Job (CPU Burst Time) First (no preemption, as we did in class)

5) Extra credit - Shortest Job (CPU Burst Time) First (with preemption, as we mentioned at the end of class)

For each simulation, if multiple processes would be placed into the ready queue at the same time, give highest priority to a newly arriving process, next a process completing its Blocked wait time (if two are unblocked at the same time, prefer the one that waited the longest), and lastly a process that has just finished its time slice.

Turn in the time line for each simulation (you may draw these by hand and scan them, or do them electronically). Also compute the turnaround time, queue waiting time, and response time for each process, as well as the percentage idle time for the CPU overall and the number of context switches. Indicate which scheduling scheme you think performed the best.
Answer #1

1. First Come First Serve

Process Arrival Time CPU Burst I/O Burst Total CPU Time Completion Time Turn Around Time(CT-AT) Waiting Time (TAT-BT) Response Time
A 0 4 4 9 4 4 0 0
B 3 2 3 7 6 3 1 1
C 6 5 1 11 11 5 0 0
D 12 1 1 5 13 1 0 0

Time line

Process A B C None D None
Time 0 4 6 11 12 13

Percentage ideal time = (1/13)x100% = 7.69%

Number of context switches = 4

2. Round robin with time slice 1:

Process Arrival Time CPU Burst I/O Burst Total CPU Time Completion Time Turn Around Time(CT-AT) Waiting Time (TAT-BT) Response Time
A 0 4 4 9 5 5 1 0
B 3 2 3 7 6 3 1 0
C 6 5 1 11 11 5 0 0
D 12 1 1 5 13 1 0 0

Time line

Process A B A B C None D None
Time 0 3 4 5 6 11 12 13

Percentage ideal time = (1/13)x100% = 7.69%

Number of context switches = 6

3. Round robin with time slice 3:

Process Arrival Time CPU Burst I/O Burst Total CPU Time Completion Time Turn Around Time(CT-AT) Waiting Time (TAT-BT) Response Time
A 0 4 4 9 6 6 2 0
B 3 2 3 7 5 2 0 0
C 6 5 1 11 11 5 0 0
D 12 1 1 5 13 1 0 0

Time line

Process A B A C C None D None
Time 0 3 5 6 9 11 12 13

Percentage ideal time = (1/13)x100% = 7.69%

Number of context switches = 5

4. Shortest job first :

Process Arrival Time CPU Burst I/O Burst Total CPU Time Completion Time Turn Around Time(CT-AT) Waiting Time (TAT-BT) Response Time
A 0 4 4 9 4 4 0 0
B 3 2 3 7 6 3 1 1
C 6 5 1 11 11 5 0 0
D 12 1 1 5 13 1 0 0

Time line

Process A B C None D None
Time 0 4 6 11 12 13

Percentage ideal time = (1/13)x100% = 7.69%

Number of context switches = 4

5. Shortest job first(with preemption) :

Process Arrival Time CPU Burst I/O Burst Total CPU Time Completion Time Turn Around Time(CT-AT) Waiting Time (TAT-BT) Response Time
A 0 4 4 9 4 4 0 0
B 3 2 3 7 6 3 1 1
C 6 5 1 11 11 5 0 0
D 12 1 1 5 13 1 0 0

Time line

Process A B C None D None
Time 0 4 6 11 12 13

Percentage ideal time = (1/13)x100% = 7.69%

Number of context switches = 4

Shortest Job first with preemption performs the best.